# SolarGPT-QA: A Domain-Adaptive Large Language Model for Educational Question Answering in Space Weather and Heliophysics

### v3

@@ IV-A Training Details and Compute
All experiments were conducted on a Linux server with dual Intel Xeon Gold 5220R CPUs (24 cores each, 2.20 GHz) and four NVIDIA RTX A5000 GPUs (24 GB VRAM). Individual training runs used a single GPU, with GPU selection controlled at runtime. To support reproducibility and transparent reporting of compute usage, our training scripts automatically log the total number of optimization steps, wall-clock training time, and peak GPU memory consumption to a training_log.json file. The total number of training tokens can be derived from these logs as:
$\text{tokens}\approx(\text{total steps})\times 512\times 8,$
corresponding to the sequence length and effective batch size used during training. All of our code used in this experiment is available on GitHub [^1] https://github.com/hlabcsusu/SolarGPT.
IV-B Dataset
@@
IV-B1 Domain Corpus for DAPT
The heliophysics corpus is curated from peer-reviewed research articles published between 2010 and 2024, corresponding to the operational era of the SDO. Articles were selected from high-impact journals with an h5-index greater than 40 to ensure scientific rigor and relevance. These venues include The Astrophysical Journal (ApJ), The Astrophysical Journal Supplement Series (ApJS), Astronomy & Astrophysics (A&A), Space Weather, and Frontiers in Astronomy and Space Sciences, among others.
Papers were filtered using eruption-related keyword co-occurrence such as solar flare, coronal mass ejection (CME), and solar energetic particles (SEP). Figure 3 shows the distribution of eruption-related research articles across five leading heliophysics journals between 2010 and 2024, stratified by individual and combined event types. This analysis identifies 1,570 relevant papers from these venues alone, and extending the selection to all qualifying journals yields an estimated 5,000–7,000 research articles used for domain adaptation. All documents used were obtained from publicly accessible sources, including open-access journal articles and author-provided PDFs available through publisher websites or institutional repositories. No proprietary databases or restricted-access collections were used. The corpus was used solely for internal model training and was not redistributed in raw or processed form, in accordance with publisher terms of use.
@@ IV-B1 Domain Corpus for DAPT
All documents were extracted from PDF format and cleaned by removing references, citations, figure captions, URLs, and formatting artifacts. Corpus-level statistics confirm strong coverage of eruption-related concepts. In particular, the word cloud shown in Fig. 4 highlights the dominance of key heliophysics terms such as solar, CME, flare, SEP, magnetic field, particle, and shock, indicating that the corpus effectively captures the semantic structure and terminology of solar eruptive phenomena.
IV-B2 QA Dataset Construction
The supervised QA dataset was constructed using a human-in-the-loop approach and consists of approximately 5,000 question–answer pairs derived from curated heliophysics literature. Large language models were used to generate initial QA candidates, guided by expert-authored few-shot exemplars that defined scientific scope and pedagogical intent. All QA pairs were subsequently reviewed, edited, and validated by human domain experts to ensure factual correctness, appropriate difficulty, and alignment with K–12 learning objectives, while filtering stylistic artifacts associated with synthetic generation.
IV-B3 Train–Test Split and Difficulty Levels
The training set consists of 30 QA pairs derived from the same corpus used during domain-adaptive pretraining. In contrast, the test set is constructed from QA pairs extracted from different PDF documents that are not included in the training corpus, ensuring document-level separation and reducing information leakage. To support a fine-grained educational evaluation, the test set is further divided into three levels of difficulty (10 Easy, 15 Medium, 5 Hard):
  • Easy: Basic factual and definitional questions, such as identifying solar phenomena or simple cause–effect relationships, suitable for elementary and middle school learners.
  • Medium: Conceptual reasoning questions that require explanation of physical processes or interactions (e.g., how CMEs interact with Earth’s magnetosphere), aligned with middle to early high school understanding.
  • Hard: Integrative questions that require synthesizing multiple concepts or explaining observational consequences (e.g., relationships between flares, CMEs, and SEPs), suitable for advanced high school students.
This tiered evaluation design enables systematic assessment of SolarGPT-QA across increasing levels of conceptual complexity while maintaining consistent scientific grounding.
@@ IV-C Baseline
We evaluate SolarGPT-QA against several strong general-purpose large language models that are widely used for scientific question answering and education. The baselines include ChatGPT (GPT-4o), Grok (Grok-1.5), Claude (Claude 3 Sonnet), DeepSeek (DeepSeek-V2), and Gemini (Gemini 1.5 Pro). All baseline models are evaluated under a zero-shot setting using identical prompts and decoding parameters, and without retrieval augmentation, to ensure a fair and controlled comparison. In addition, we include an instruction-tuned variant of ChatGPT, where the model is prompt-tuned using a small number of representative examples (few-shot setting), in order to compare SolarGPT-QA against a highly optimized general-purpose model with explicit pedagogical alignment.
IV-D Evaluation Method
We conducted a human pairwise evaluation on 30 test questions, stratified by difficulty (10 Easy, 15 Medium, 5 Hard). For each question, model outputs were independently evaluated by two space weather domain experts, resulting in 60 total judgments per model comparison. Annotators were presented with anonymized and randomly ordered responses and followed a standardized evaluation rubric assessing (i) factual correctness, (ii) clarity of explanation, (iii) age-appropriateness, and (iv) overall educational quality. For each question, annotators assigned one of three outcomes: Win, Tie, or Loss relative to the baseline model. Inter-annotator agreement was measured using Cohen’s $\kappa$, yielding $\kappa=0.72$, which indicates substantial agreement. The win rate is defined as the proportion of questions where model is judged to outperform the baseline. Given the limited number of annotators, the results are indicative rather than statistically conclusive.
IV-E Results
We evaluate model performance using human pairwise comparison to assess response quality in educational question answering for space weather and heliophysics. Figure 5 presents the overall win-rate comparison across multiple baselines.
ChatGPT (instruction-tuned, 10-shot) achieves the highest win rate among all evaluated models, indicating that human evaluators consistently prefer its responses. This result highlights the substantial benefit of instruction tuning and few-shot prompting for improving clarity and response relevance. SolarGPT-QA performs competitively and outperforms all other zero-shot and general-purpose baselines, but remains inferior to ChatGPT when enhanced with instructional alignment. ChatGPT (zero-shot) and Grok demonstrate moderate performance, while Claude, DeepSeek, and Gemini obtain lower win rates, underscoring the challenges faced by general-purpose models in specialized educational contexts.
Figure 5: Human evaluation win-rate comparison across models.
To further assess practical educational effectiveness, we conduct a small pilot student comprehension study involving 15 students. Each student was asked to read the answers generated by each model and report whether they understood the response. Figure 6 shows the number of students who reported understanding the answers produced by each model.
ChatGPT (instruction-tuned, 10-shot) achieves the highest level of student comprehension, with all 15 students reporting understanding its responses. SolarGPT-QA follows closely, with 12 students indicating comprehension. ChatGPT (zero-shot) and Grok demonstrate moderate comprehension levels, while Claude, Gemini, and DeepSeek show lower student understanding overall. These results further suggest that instructional alignment and pedagogical optimization play a crucial role in educational question answering, beyond domain specialization alone.
Figure 6: Student comprehension comparison across models.
IV-F Ablation Study
@@
We conduct an ablation study to analyze the impact of domain-adaptive pretraining (DAPT) and educational fine-tuning in SolarGPT-QA. We compare four model variants: the base LLaMA-3 model, LLaMA-3 with DAPT only, LLaMA-3 with QA fine-tuning only, and the full SolarGPT-QA model. DAPT significantly improves scientific grounding and correct use of heliophysics terminology, but often produces explanations that are too technical for K–12 learners. QA fine-tuning alone improves readability and storytelling but lacks sufficient domain depth, leading to shallow or incomplete explanations. Table I shows SolarGPT-QA achieves the highest win rate (75%) and student comprehension (80%), suggesting that combining domain-adaptive pretraining with educational fine-tuning yields the strongest performance.
TABLE I: Ablation Study Results for SolarGPT-QA
Model Variant Win Rate (%) Student Understanding (%)
Base LLaMA-3-8B 28 20 (3/15)
LLaMA-3 + DAPT 54 40 (6/15)
LLaMA-3 + QA Fine-Tuning 48 67 (10/15)
SolarGPT-QA 75 80 (12/15)
The results indicate that SolarGPT-QA benefits significantly from domain-adaptive pretraining and fine-tuning. While ChatGPT and Grok often generate factually correct responses, they frequently use advanced terminology or omit explanatory details, which reduces their educational effectiveness.
V Discussion and Conclusion
V-A Discussion
The results presented in this work demonstrate that domain adaptation alone is insufficient for educational effectiveness and that pedagogical alignment is a critical and often overlooked dimension when deploying large language models for science education. While general-purpose LLMs such as ChatGPT and Grok exhibit strong linguistic fluency and factual recall, our findings show that they frequently fail to deliver explanations that are accessible, coherent, and age-appropriate for K–12 learners in the context of space weather and heliophysics.
In the human pairwise evaluation, instruction-tuned ChatGPT (10-shot) achieves the highest overall win rate, indicating that evaluators consistently favor responses that are explicitly aligned with instructional intent and teaching style. Despite this, SolarGPT-QA performs competitively and outperforms all zero-shot general-purpose baselines, achieving a 75% overall win rate. This result suggests that domain-adaptive pretraining combined with educational fine-tuning can substantially narrow the gap between zero-shot domain models and heavily instruction-tuned general LLMs, without relying on in-context examples at inference time.
The student comprehension study further refines this interpretation. While instruction-tuned ChatGPT achieves perfect comprehension among participants, SolarGPT-QA follows closely, with a large majority of students reporting clear understanding of its explanations. In contrast, comprehension levels for zero-shot general-purpose models remain noticeably lower. These results indicate that SolarGPT-QA succeeds in translating complex heliophysics concepts into accessible explanations, even if it does not surpass the strongest instruction-aligned baseline.
The ablation study provides critical insight into why this performance emerges. Domain-adaptive pretraining (DAPT) alone improves scientific grounding but tends to preserve expert-level language, while QA fine-tuning alone improves readability at the expense of domain depth. The full SolarGPT-QA model achieves the strongest balance between these objectives, producing explanations that are both scientifically grounded and pedagogically accessible. Importantly, SolarGPT-QA reaches this balance without external prompting strategies, highlighting the effectiveness of integrating educational objectives directly into the training process.
These findings have broader implications beyond heliophysics. They suggest that many existing scientific LLMs, while impressive for expert-level analysis, are poorly suited for educational deployment without explicit adaptation. As AI systems increasingly enter classrooms, museums, and public outreach platforms, models that fail to account for learner expertise risk reinforcing misunderstanding or disengagement. SolarGPT-QA provides preliminary evidence that targeted domain adaptation combined with educational fine-tuning can effectively bridge this gap.
V-B Limitations
Both training and evaluation QA pairs were initially generated using large language models. Although we enforce strict document-level separation between train and test sources and apply expert human review to reduce synthetic artifacts, residual stylistic overlap or generator-specific priors may inflate performance. Some baseline models share vendors with the QA generators, potentially introducing bias due to shared pretraining or alignment conventions. We partially mitigate this by evaluating multiple independent model families under identical zero-shot settings and by using ablation studies to isolate the effects of domain adaptation and educational fine-tuning. Stronger validation using fully human-authored test sets is left to future work.
V-C Future work
The present study represents an initial, proof-of-concept exploration of domain-adapted large language models for educational question answering in space weather and heliophysics. Several important extensions are planned as part of the broader SolarGPT research agenda.
First, future work will focus on developing SolarGPT-Core, a large-scale domain-adapted foundation model trained on an expanded corpus of peer-reviewed heliophysics literature. Scaling to larger architectures and longer context windows is expected to improve scientific depth, reasoning consistency, and coverage of rare but high-impact solar events.
Second, we plan to introduce SolarGPT-PredEx, a multimodal prediction-and-explanation model that integrates time-series measurements and solar imagery with language modeling. Unlike the current SolarGPT-QA system, SolarGPT-PredEx will directly reason over observational data to produce both event predictions and data-grounded natural language explanations, enabling more transparent and interpretable space weather forecasting.
Third, future versions of SolarGPT-QA will incorporate retrieval-augmented generation (RAG), curriculum-aware prompting, and adaptive difficulty control to support classroom deployment and personalized learning. Larger and more diverse human evaluation studies, particularly involving teachers and students across multiple grade levels, will be conducted to rigorously assess educational effectiveness.
Finally, all components of the SolarGPT ecosystem will be integrated into a unified research and outreach platform that bridges prediction, explanation, and education, supporting both scientific discovery and public understanding of heliophysics. The current work establishes the feasibility of this vision while clearly delineating its preliminary scope.
V-D Conclusion
In this work, we introduced SolarGPT-QA, a domain-adapted large language model designed for educational question answering in space weather and heliophysics. By combining domain-adaptive pretraining on peer-reviewed heliophysics literature with pedagogically aligned question–answer fine-tuning, SolarGPT-QA aims to bridge expert scientific knowledge and accessible K–12 instruction.
Human pairwise evaluation results show that while instruction-tuned ChatGPT (10-shot) achieves the strongest overall performance, SolarGPT-QA outperforms all zero-shot general-purpose baselines, and suggests strong student comprehension and explanatory clarity. While these findings are preliminary and based on limited sample sizes, they provide initial evidence that integrating domain adaptation with educational fine-tuning can enhance the clarity and accessibility of scientific explanations. Ablation experiments further demonstrate that neither domain-adaptive pretraining nor educational fine-tuning alone is sufficient; rather, their combination is important for balancing scientific accuracy with pedagogical effectiveness.
Beyond its direct use, this work shows a clear method for building educational science language models. It highlights that good AI learning tools should include teaching goals, not just subject knowledge. SolarGPT-QA is a first step toward a larger SolarGPT system that brings together prediction, explanation, and education in heliophysics.
VI Acknowledgements

### v4

@@ 4.1 Training Details and Compute
We conducted both DAPT and QA fine-tuning on Meta-LLaMA-3-8B using Low-Rank Adaptation (LoRA). To reduce memory usage, the pretrained LLaMA-3-8B backbone was loaded in 8-bit precision, while training computations for the LoRA parameters were performed in FP16 (16-bit floating point). Rather than updating all model parameters, LoRA adapters were inserted into the query and value projection layers of the attention modules (i.e., q_proj and v_proj), with rank r=16r{=}16, scaling $\alpha{=}32$, and dropout p=0.05p{=}0.05. Training used the AdamW optimizer with a learning rate of $2\times 10^{-4}$, linear warmup of 100 steps, and a maximum sequence length of 512 tokens. The per-device batch size was 1 with gradient accumulation of 8, resulting in an effective batch size of 8 sequences per optimization step. Both DAPT and QA fine-tuning were run for 3 epochs. All experiments were conducted on a Linux server with dual Intel Xeon Gold 5220R CPUs (24 cores each, 2.20 GHz) and four NVIDIA RTX A5000 GPUs (24 GB VRAM). Individual training runs used a single GPU, with GPU selection controlled at runtime. To support reproducibility and transparent reporting of compute usage, our training scripts automatically log the total number of optimization steps, wall-clock training time, and peak GPU memory consumption to a training_log.json file. The total number of training tokens can be derived from these logs as: $\text{tokens}\approx(\text{total steps})\times 512\times 8$, corresponding to the sequence length and effective batch size used during training. All of our codes and datasets used in this experiment are available on GitHub [^1] https://github.com/hlabcsusu/SolarGPT.
4.2 Dataset
@@
4.2.1 Domain Corpus for DAPT
The heliophysics corpus is curated from peer-reviewed research articles published between 2010 and 2024, corresponding to the operational era of the SDO. Articles were selected from high-impact journals with an h5-index greater than 40 to ensure scientific rigor and relevance. These venues include The Astrophysical Journal (ApJ), The Astrophysical Journal Supplement Series (ApJS), Astronomy & Astrophysics (A&A), Space Weather, and Frontiers in Astronomy and Space Sciences, among others. Papers were filtered using eruption-related keyword co-occurrence such as solar flare, coronal mass ejection (CME), and solar energetic particles (SEP). Fig. 3 shows the distribution of eruption-related research articles across five leading heliophysics journals between 2010 and 2024, stratified by individual and combined event types. This analysis identifies 1,570 relevant papers from these venues alone, and extending the selection to all qualifying journals yields an estimated 5,966 research articles used for domain adaptation. All documents used were obtained from publicly accessible sources, including open-access journal articles and author-provided PDFs available through publisher websites or institutional repositories. No proprietary databases or restricted-access collections were used. The corpus was used solely for internal model training and was not redistributed in raw or processed form, in accordance with publisher terms of use. All documents were extracted from PDF format and cleaned by removing references, citations, figure captions, URLs, and formatting artifacts. Corpus-level statistics confirm strong coverage of eruption-related concepts. In particular, the word cloud shown in Fig. 4 highlights the dominance of key heliophysics terms such as solar, CME, flare, SEP, magnetic field, particle, and shock, indicating that the corpus effectively captures the semantic structure and terminology of solar eruptive phenomena.
@@ 4.2.1 Domain Corpus for DAPT
4.2.2 QA Dataset Construction
The supervised QA dataset consists of approximately 5,000 question–answer pairs derived from curated heliophysics literature using a human-in-the-loop approach, where initial LLM-generated candidates were rigorously validated by domain experts for scientific accuracy and K–12 pedagogical alignment. To ensure a robust assessment, a separate test set of 300 QA pairs was extracted from PDF documents that were not included in the training corpus, effectively eliminating information leakage. This test set is stratified into three difficulty levels: easy questions cover basic definitions and solar phenomena for elementary learners; medium questions require reasoning about physical processes like CME interactions with the magnetosphere; and hard questions demand the synthesis of complex concepts, such as the relationship between flares and SEPs. This tiered structure enables a systematic evaluation of SolarGPT-QA across varying levels of conceptual depth and learner maturity while maintaining consistent scientific grounding.
@@ 4.3 Baseline
We evaluate SolarGPT-QA against several strong general-purpose large language models that are widely used for scientific question answering and education. The baselines include ChatGPT (GPT-4o), Grok (Grok-1.5), Claude (Claude 3 Sonnet), DeepSeek (DeepSeek-V2), and Gemini (Gemini 1.5 Pro). All baseline models are evaluated under a zero-shot setting using identical prompts and decoding parameters without retrieval augmentation, to ensure a fair and controlled comparison. In addition, we include an instruction-tuned variant of ChatGPT, where the model is prompt-tuned using a small number of representative examples (few-shot setting), in order to compare SolarGPT-QA against a highly optimized general-purpose model with explicit pedagogical alignment.
4.4 Evaluation Method
To evaluate our model, we employ an LLM-as-a-Judge [gu2024survey] framework, utilizing a GPT-4o to conduct automated pairwise assessments in place of manual human review. We test model responses against a stratified set of 300 questions consisting of 100 easy, 150 medium, and 50 hard test instances in which outputs from SolarGPT-QA and a baseline model are presented to the evaluator in a randomized, anonymized order to eliminate positional bias. The judge assesses each pair based on a structured rubric tailored for K–12 heliophysics education, specifically measuring factual correctness, clarity of explanation, age appropriateness, and educational usefulness. Each comparison results in a win, tie, or loss for SolarGPT-QA, with the final win rate calculated as the proportion of preferred responses. We repeat each comparison multiple times and the final win rate is reported as the mean across runs with standard deviation.
4.5 Results
We evaluate SolarGPT-QA against several strong general-purpose large language models using LLM-based pairwise comparison. Fig. 5 presents the win-rate comparison across models. ChatGPT (instruction-tuned, 10-shot) achieves the highest win rate at 83%, indicating that the judge model consistently prefers its responses. SolarGPT-QA achieves a win rate of 72%, outperforming all zero-shot baselines, including ChatGPT (55%), Grok (50%), Claude (45%), Gemini (42%), and DeepSeek (38%). These results suggest that domain-adaptive pretraining combined with educational fine-tuning substantially improves response quality in specialized scientific question answering, while remaining competitive with strongly prompted general-purpose models.
Figure 5: LLM-as-a-Judge win-rate comparison across models. Bars represent the mean win rate over repeated evaluation runs, and error bars denote the standard deviation. Numerical labels above each bar indicate the mean win rate.
4.6 Ablation Study
@@
We conduct an ablation study to analyze the impact of domain-adaptive pretraining (DAPT) and educational QA fine-tuning in SolarGPT-QA. Four model variants are compared: the base LLaMA-3 model, LLaMA-3 with DAPT only, LLaMA-3 with QA fine-tuning only, and the full SolarGPT-QA model. DAPT significantly improves scientific grounding and the correct use of heliophysics terminology. However, responses produced by the DAPT-only model often contain explanations that are overly technical for K–12 learners. In contrast, QA fine-tuning alone improves readability and explanatory structure but sometimes lacks sufficient domain depth, leading to incomplete or simplified explanations. Table 1 shows that the full SolarGPT-QA model achieves the highest win rate, indicating that combining domain-adaptive pretraining with educational fine-tuning yields the strongest overall performance. The ablation results indicate that both components contribute complementary benefits. Domain-adaptive pretraining strengthens scientific accuracy and terminology usage, while QA fine-tuning improves explanation quality and pedagogical clarity. Their combination produces responses that better balance scientific rigor with educational accessibility.
Table 1: Ablation study evaluating DAPT and QA fine-tuning.
Model Variant Win Rate (%, mean $\pm$ std)
Base LLaMA-3-8B $28.0\pm 1.9$
LLaMA-3 + DAPT $54.0\pm 2.4$
LLaMA-3 + QA Fine-Tuning $48.0\pm 2.2$
SolarGPT-QA $\mathbf{72.0\pm 3.2}$
5 Discussion and Conclusion
5.1 Discussion
The results demonstrate that domain adaptation alone is insufficient for effective educational explanations and that pedagogical alignment is an important factor when deploying large language models for science education. While general-purpose LLMs such as ChatGPT and Grok exhibit strong linguistic fluency and factual recall, they often produce explanations that are either overly technical or insufficiently structured for K–12 learners in specialized domains such as heliophysics. Using an LLM-as-a-Judge evaluation framework, we find that instruction-tuned ChatGPT (10-shot) achieves the highest overall win rate, indicating strong performance when explicit instructional alignment is introduced through prompting. SolarGPT-QA performs competitively and outperforms all zero-shot general-purpose baselines, achieving a 72% win rate. This suggests that combining domain-adaptive pretraining with educational fine-tuning can significantly improve the quality of explanations in specialized scientific domains without requiring few-shot prompting during inference. The ablation study further clarifies the contribution of each training stage. Domain-adaptive pretraining improves scientific grounding and terminology usage but tends to preserve expert-level language. QA fine-tuning improves readability and explanatory structure but may lack domain depth when used alone. The full SolarGPT-QA model offers the best balance, giving explanations that are both scientifically accurate and easy to understand. More broadly, these findings suggest that many existing scientific LLMs are optimized for expert-level reasoning rather than educational communication. As AI systems increasingly support science outreach and education, integrating domain knowledge with pedagogical alignment will be essential for producing explanations that are both accurate and accessible.
5.2 Limitations
Both training and evaluation QA pairs were initially generated using large language models. Although strict document-level separation between training and test sources is enforced, residual stylistic artifacts from synthetic generation may remain. Additionally, the evaluation relies on an LLM-as-a-Judge framework, which may introduce bias associated with the judge model’s own preferences or training distribution. While pairwise evaluation and randomized response ordering help mitigate these effects, further validation using independent evaluation methods remains important for future work.
5.3 Future Work
This study represents an initial step toward domain-adapted language models for heliophysics education. Future work will focus on developing SolarGPT-Core, a larger domain-adapted foundation model trained on an expanded corpus of heliophysics literature to improve scientific depth and coverage of rare solar events. We also plan to develop SolarGPT-PredEx, a multimodal model that integrates solar imagery and time-series measurements with language modeling to generate both space weather predictions and natural language explanations grounded in observational data. Additional improvements to SolarGPT-QA will explore retrieval-augmented generation (RAG), curriculum-aware prompting, and adaptive difficulty control to support classroom and outreach applications.
5.4 Conclusion
We introduced SolarGPT-QA, a domain-adapted large language model designed for educational question answering in space weather and heliophysics. By combining domain-adaptive pretraining on heliophysics literature with QA fine-tuning, which improves the clarity and accessibility of scientific explanations. Evaluation using an LLM-as-a-Judge framework shows that SolarGPT-QA outperforms multiple zero-shot general-purpose models and approaches the performance of strongly instruction-tuned baselines. Ablation studies further demonstrate that combining domain adaptation with educational fine-tuning is important for balancing scientific accuracy with pedagogical clarity. Overall, this work provides a practical approach for building domain-specific educational language models and represents an initial step toward a broader SolarGPT ecosystem that integrates scientific knowledge, explanation, and education in heliophysics.
{credits}
5.4.1 Acknowledgements